October 23, 2001
Report by Bob Feldman
2:04 GMT Alvin submerged.
3:30 GMT Bottom sighted, glassy basalt off axis.
3:45 GMT ALS visually sighted, hose inspected, found to be unkinked, we spotted a crab on the release and attempted to catch it. It turned out to be a galatheid so we stopped and proceeded to release the ALS.
4:00 GMT On ALS preparing to release, X=4497, Y=77994.
4:17 GMT ALS released, Anthony first grabbed the bungee then released and grabbed the yellow loop. Both Tony and Bruce expressed polite dissatisfaction with the bungee mechanism and a desire to assist in future science-free vehicle design. Sample page filled out. Left for TICA site.
4:35 GMT Arrived TICA, visually interrogated several Riftia fields.
4:44 GMT Sampled Riftia field 1.
4:59 GMT Sampled Riftia field 2.
5:22 GMT Sampled Riftia field 3. Sample page filled out. Moved to Bio9.
5:53 GMT Arrive Bio9, surveyed Alvinella colonies.
5:59 GMT HiT probe, 81.3C but no Alvinellids.
6:16 GMT HiT probe temps up to 65C but no Alvinellids
6:19 GMT Spotted plate marker 18, Bruce and Tony were not sure exactly where we were.
6:23 GMT Spotted Dave, verifying Bio9 position.
6:38 GMT Moved to X=4621, Y=78020, HiT temp at 66°C
6:42 GMT Sampled Alvinella clump at X=4816.016, Y=78020.286, Heading 41,44. HiT temp was 66°C, made 2 grabs, all animals put in small biobox. No Sipper data. Sample sheet.
6:54 GMT Sipper samples taken at X=4816.671, Y=78019.421, Hdg 320. HiT temp=82°C, fired Sippers 12, 1, 2 but pump not on. Sipper sample 3 is ok. Temp was 102°C. Sample sheet filled out.
7:42 GMT Sipper sample 5 fired (4 was misfire). HiT =133°C, x=4618.80, y=78023.071. Hdg = 342.
7:51 GMT Alvinellid samples from same place as Sippers. HiT 82°-133°C. x=4618.88, y=78023.071.
8:08 GMT Left Bio9 heading west out of ASC for crabs and basalt.
8:20 GMT Sipper 6 fired at base of glassy basalt, temp = 1.8°C, x= 4573.338, y=77989.359.
8:48 GMT John Holloway basalt taken at x=4561.578, y=77989.961 put in milk crate, temp =1.9°C
8:59 GMT Basalt sample taken at 45661.574, y=77989.516, temp=1.8°C, 3 total, 1 small, 1 med, 1 large.
9:00 GMT Attempted to catch crabs.
9:14 GMT Dropped weights.